OpenStudy (anonymous):

choose the correct simplification of a^5b^6/ a^4b^4

OpenStudy (anonymous):

a b^2

OpenStudy (anonymous):

how did you get the answer?

OpenStudy (anonymous):

\[\frac{ a^5 b^6}{ a^4b^4 } = ab^2. \] When you divide, you subtract the exponents

OpenStudy (anonymous):

5-4 = 1 = a^1 6-4 = 2 = b^2

OpenStudy (anonymous):

choose the correct simplification of (x^3/y^5)^2

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

Power to a power, multiply the exponents. x^(3*2) y^(5*2)

OpenStudy (anonymous):

x^6y^10?

OpenStudy (anonymous):

Is it (x6y)^10, or (x^6y)^(10)?

OpenStudy (anonymous):

it would be x^6 over y^10